an unknown gas effuses 3.87 times faster than H2. What is the molar mass of the gas?
Answer #1

we have a relation

R1 /R2 = (M2/M1)

R1 = 1.0

R2 = 3.87

M1 = 2 g /mol (for H2)

M2 = ?

1.0 / 3.87 = M2 / 2 square on both sides

1.0 / 14.9769 = M2 / 2.0

0.0668 = M2 / 2.0

M2 = 0.133 g /mol

answer = 0.133 g /mol